OpenStudy (anonymous):

I need some help i have tried to do this but i keep on doing something wrong! Best explanation I can give. The data can be modeled by the function f(x)= 8 + 38ln x, where f(x) is the percentage of new cellphones with cameras x years after 2002. 2003-8% 2004-34% 2005-50% 2006-63% 2007-68% Thank you much! A)use the function to determine the percentage of new cellphones with cameras in 2006. Round to the nearest whole percent.Does this overestimate or underestimate the percent by the graph? By how much? Another part couldn't fit, listed below

OpenStudy (anonymous):

B) If trends show from 2003 through 2007 continue, use the function to determine by which year all new cellphones will be equipped with cameras

OpenStudy (anonymous):

So in A) your x value is going to be 2006-2002=4, so plugging f(4) in, you get 8+38ln(4)=60.679.... thus it over estimates the data given by 0.679....%

OpenStudy (anonymous):

b) is asking you to solve the equation for "100%" which is 100=8+38ln(x), can you do this?

OpenStudy (anonymous):

I believe I can. Thank you :-)